# Note on KSW-allowability of Wine-Glass saddles

### 24 Mar 2026 v1

@@ Abstract
In this note we consider no-boundary instantons and wine-glass geometries which are of interest in the context of quantum cosmology. While the former appears as a dominant saddle in the path-integral, the later sub-dominant saddle has been argued to have a longer inflationary phase of the Universe. Kontsevich-Segal-Witten (KSW)-allowability criterion which classifies geometries on the basis of the requirement of having a meaningful QFT on it, pushes one to analyse the allowability of the saddles. In this note we do a simple study to seek answer to the allowabilty of no-boundary instantons and wine-glass geometries which are of interest in quantum cosmology. Our simple analysis which make use of a milder version of KSW allowability criterion shows that no-boundary instanton is KSW allowed while wine-glass geometries are KSW disallowed.
@@ I Introduction
Understanding of gravitational physics has been significantly improved due to various studies involving Gravitational path-integrals, where they acquire center stage in investigating issues dealing with either holography or black hole physics or cosmologies. While a clean definition of path-integral in ordinary non-gravitational QFT can be arranged, such a luxury is not achievable in gravitational systems which suffers from perturbative non-renormalizability, suggesting an effective treatment of the problem tHooft:1974toh ; Deser:1974nb ; Deser:1974cz ; Goroff:1985sz ; Goroff:1985th ; vandeVen:1991gw . Beside this, the choice of contour enabling flat spacetime QFT path-integral convergent is not available. On other hand, the conformal mode leading to unboundedness of the Euclidean gravitational path-integral is suggestive of working with the gravitational path-integral directly in Lorentzian signature which however is highly oscillatory Gibbons:1978ac . As a result the path-integral over the geometries becomes a challenging problem which is sensitive to boundary choices and contours in the space of all possible complex metrics.
@@
In a semi-classical approximation, the gravitational path-integral reduces to summation over contributions coming from various saddle geometries. Typically for given boundary choices, multiple saddle-geometries (which can be complex) appear in path-integral competing with each other. Dominance of one over the other is dictated by the external inputs, where the saddles can loose their dominance as external parameters are varied. This is analogous to phase transitions occurring in non-gravitational settings. Understanding dominance of saddles is therefore crucial in gaining a deeper insight to various phases of the gravitational systems.
@@
However, when path-integral over metrics is extended to complex space, a subtlety arises. Not all such metrics are allowable, where quantum field theories can be meaningfully and sensibly defined Witten:2021nzp . Kontsevich-Segal-Witten (KSW)-allowability criterion offers a diagnostic tool to identify such complex metrics. This criterion directly arises from the requirements of convergence and well-behavedness of the path-integral, bifurcating the space of complex geometries in to allowable and non-allowable ones. A crucial question then arises regarding the allowability of various complex saddles in the path-integral, asking piercing questions regarding the region where they lie. This is however sensitive to choice of external parameters for example boundary configurations, in the sense that saddle allowability is affected by boundary choices.
@@
In the context of quantum cosmology, it is seen that different boundary choices at initial time lead to a very different Universe in the future. By demanding a Universe to be devoid of an initial singularity proposes a Universe starting from zero size (No-boundary Universe). In a simple mini-superspace model it is achievable by either fixing initial size of Universe (Dirichlet BC) or by fixing initial momentum (Neumann BC) DiTucci:2019bui ; Narain:2021bff ; Lehners:2021jmv ; DiTucci:2020weq ; Narain:2022msz ; Ailiga:2023wzl ; Ailiga:2024mmt . In the former multiple saddles compete for dominance in various regimes given by final size of Universe. The complex saddles which eventually becomes relevant in the path-integral computed via Picard-Lefschetz methods turn out to have unstable perturbations Feldbrugge:2017kzv ; DiTucci:2018fdg ; Feldbrugge:2017fcc . In the Neumann case, the complex saddles corresponds to Hartle-Hawking (HH) no-boundary geometries, that leads to stable perturbations Lehners:2021jmv ; Ailiga:2024wdx . The saddle corresponding to yje tunneling proposal is not only sub-dominant but also has unstable perturbations. Despite the appealing features, the HH-saddle however, gives rise to great puzzles: it predicts an empty Universe with least possible number of inflationary e-folds and predicts overall spatial curvature of the Universe that is significantly larger than the observed one Maldacena:2024uhs .
@@
Recently, it has been noticed that saddle configurations corresponding to wine-glass geometries eases this tension Betzios:2024oli ; Betzios:2024zhf ; Betzios:2026rbv , though such saddles don’t enjoy the privilege of being dominant in the path-integral as noticed in Lavrelashvili:2026zsw . These are wormhole saddles where the Euclidean wormhole bridges the two asymptotic regions, and are intrinsically non-perturbative configurations following from equation of motion. Such Euclidean wormhole saddles which are supported by Axion-field, can be analytically continued to Lorentzian signature, where they lead to cosmological evolution. A particular feature of these wormhole geometries is that they are asymptotically AdS in the far past. These gravitational saddles are noticed to overcome the various issues affecting the Hartle-Hawking no-boundary saddle i.e. supports long lasting inflationary period. Though it is still not well-established how from a single boundary configuration both wineglass saddles and no-boundary instanton can both arise, where the wineglass saddle is dominant.
@@
In this short note we check the KSW allowability of the wineglass saddle. This is important because if this saddle is dominant for some boundary configurations and is KSW-allowed then it becomes a favourable proposal over the Hartle-Hawking no-boundary instanton as it overcomes various issues by giving preference to longer inflationary phase. On the other hand if it is not KSW-allowed, which may very well happen, as has been seen to be the case for bouncing cosmologies or typical wormholes geometries, then it is an unphysical geometry on which sensible QFTs cannot be defined. It is therefore crucial to address this issue. In this short note we do a semi-rigorous study to analyse the KSW-allowability of the wineglass saddle.
The rest of the paper is organized as follows: In Sec. (II), we briefly review the KSW criterion. Sec. (III), details about the techniques used to check the admissibility complex geometry and particularly applied for No-boundary instanton, while Sec. (IV) deals with the admissibility of wine-glass instanton. Finally, Sec. (V) presents summary of our findings and conclusions of the paper.
@@ IV Allowability of Wine-Glass geometry
After having studied the admissibility of the NB instanton, we now proceed to study another interesting complex geometry - Wineglass wormholes with asymptotically Euclidean AdS in the far past, see figure (1(b)). Recent studies Betzios:2024oli ; Betzios:2024zhf ; Betzios:2026rbv ; Lavrelashvili:2026zsw , suggest that imposing the wine-glass wormholes as the initial condition can lead to a prolonged period of inflation, potentially resolving some of the shortcomings of the no-boundary proposal. This proposal has also been extended to include wormhole solutions supported by magnetic or axionic charges Jonas:2023ipa ; Betzios:2024oli ; Betzios:2024zhf ; Betzios:2026rbv ; Lavrelashvili:2026zsw . In the present work, however, we restrict our attention to pure gravity and follow the time contour approach presented in Hertog:2011ky . In Hertog:2011ky , utilizing the apporpriate time contour, the authors showed that in the context of AdS/CFT, the no-boundary saddle admits the a representation of regular Euclidean AdS domain wall that makes a smooth transition to a asymptotically deSitter. It is emphasized that the transition between AdS and dS regulates the volume divergences of AdS action and accounts for the phases that explain in the late time deSitter.
@@
Figure 3: A visualization of wineglass geometry and the corresponding time contour. In fig (a): the light-orange correspond to Euclidean AdS, the gray-region is the wormhole joining EAdS to EdS, the blue-region is the section of 44-sphere which is the Euclidean deSitter, the white-region is the Lorentzian deSitter. Fig (b) correspond to the time-contour: the vertical red line correspond to EAdS, the circular region depicts the wormhole with throat, the horizontal black line correspond to section of spherical geometry, while the vertical blue line correspond to Lorenztian dS.
@@
This is a negative signature representation of the geometry of Euclidean anti-deSitter space, with a cosmological constant $\bar{\Lambda}=-\Lambda$. The recent studies Betzios:2024oli ; Betzios:2024zhf ; Betzios:2026rbv shows contributions from this contour will give higher weighting to long lasting inflationary period. Now applying KSW criterion to the metric along this contour, we find
@@
which clearly shows the violation of KSW condition. This hints atleast in case of pure gravity the time contour realization of the wine-glass geometry is KSW disallowed, while NB saddle is KSW allowed. A better understanding can be gained by analyzing the $q=0$ admissibility condition i.e $\mathcal{R}e(\sqrt{g})>0$, along a complex circular contour (cyan contour in fig. (3)) that glues the Euclidean AdS geometry with Euclidean sphere. Along the circular contour, $\tau=Re^{i\theta}$, with $\theta\in[-\pi/2,0]$ the geometry given in eq. (7) reads as:
@@
From the plot (3), it is clearly evident to see the violation of $q=0$ admissibility condition for some $\theta$ interval. This supports the argument that wine-glass geometry in pure gravity with cosmological constant is KSW disallowed. Interesting, in the vanishing throat limit i.e. $R\to 0$ one achieves no-boundary instanton, which is KSW - allowed. Furthermore, for better and conclusive insight, we perform the Ridge criterion analysis. The Ridge criterion for wine-glass is illustrated in the figure (3). The analysis reveals that the wine-glass geometry — moreover, any asymptotically EAdS configuration — violates the KSW criterion. This is because, as illustrated in Fig.(3), the asymptotically EAdS region is surrounded by the disallowed region, implying that there exists no allowable contour connecting an asymptotically EAdS geometry to an asymptotically de Sitter configuration. This hints atleast from the time contour representation point of view, the wine-glass instanton is KSW disallowed, while NB saddle is KSW allowed.
@@
In this short paper we consider class of saddle geometries which resemble like wine-glass and have recently gained attention in the area of quantum cosmology due to their ability to lead to a longer phase of inflation thereby overcoming the issues affecting the Hartle-Hawking no-boundary instanton. These wine-glass saddles are wormholes which typically connect the Euclidean AdS in the far past to a section of Euclidean $S^{4}$ which glues to Lorentzian deSitter at the equator (see figure 1(b)). While these saddles have been shown to be sub-dominant compared to No-boundary instanton Lavrelashvili:2026zsw , they still come under same family as the No-boundary instanton emerges from wine-glass in the limit of vanishing throat size (when pinching occurs).
@@
We then apply these tools to the wineglass-wormhole geometries. The time contour leading to this class of geometries is shown in figure (3). Geometries with negligible throat size correspond to contours involving sharp turns at the origin of the complex $\tau$-plane, which is singular ($R\to 0$). Non-zero throat size involve contours starting at $\tau_{y}=$-$\infty$, where the geometry is EAdS. It stops at $\tau_{y}=-\tau^{(0)}_{y}<0$, leading to a non-zero throat size. The contour then ventures in complex plane to connect to No-boundary instanton contour starting at $\tau_{x}=-\tau^{(0)}_{x}>0$. The transition from EAdS to EdS happens entirely via complex geometries.
@@
Points lying in the far past correspond to EAdS geometries. This mentioned in eq. (17) shows that all entries are negative, immediately indicating its KSW disallowability. While this is not conclusive enough, we make use of KSW allowability criterion for $q=0$-form fields. This correspond to checking ${\rm Re}(\sqrt{g})>0$ for various points on the complex plane, thereby dividing the whole complex-$\tau$ plane into two parts. This is shown in figure (a) of 4 (green region is allowed while light-orange region is disallowed). This is a bit crude test but gives sufficient hints. It can be clearly from figure (a) of 4 that the wormhole (cyan-line) lies partially in allowed and partially disallowed region. This indicates that KSW-allowability following from 0-form field is compromised.
@@
To strengthen our analysis, we check the allowability by using ridge criterion. As this test is comparatively more definitive, it offers further assurance of KSW-disallowability of class of wineglass geometries. This test shows (see figure (b) in 4) that while contours corresponding to Euclidean (black-line) and Lorentzian (blue-line) deSitter are lying in KSW allowed region, the contour (red-line) corresponding to EAdS lies in KSW-disallowed region. The contour corresponding to the wormhole (cyan-line) lies partially in allowed and partially in disallowed region. This indicates that contour corresponding to wineglass is KSW-disallowed.
@@
One must however add that KSW-allowability of saddles depend on quite a few things and can get affected depending on boundary choices, presence of additional fields and their interactions. The present study doesn’t take those into consideration. Here we just checked that given a wineglass-geometry, is it KSW-allowed? A more rigorous study involving extremal curves along with presence of additional interacting fields is missing in the current work and we hope to address those in the future investigations.

### 30 Mar 2026 v2

# Note on KSW-allowability of Wine-Glass Geometry

@@ Abstract
In this note we consider no-boundary instantons and wine-glass geometries which are of interest in the context of quantum cosmology. While the former usually appears as a dominant saddle in the path-integral, the wineglass geometry can become dominant saddle in some situations. The later has been argued to have a longer inflationary phase of the Universe. Kontsevich-Segal-Witten (KSW)-allowability criterion which classifies geometries on the basis of the requirement of having a meaningful QFT on it, pushes one to analyse the allowability of the various geometries. In this note we do a simple study to seek answer to the allowabilty of no-boundary instantons and wine-glass geometries, where the later is obtained via analytical continuation of Lorentzian deSitter in pure gravity. Our simple analysis which make use of a milder version of KSW allowability criterion shows that no-boundary instanton is KSW allowed while wine-glass geometries obtained via such analytic continuation in pure-gravity are KSW disallowed. This study however doesn’t covers wineglass saddles arising in gravity coupled with matter theories.
@@ I Introduction
Understanding of gravitational physics has been significantly improved due to various studies involving Gravitational path-integrals, where they acquire center stage in investigating issues dealing with either holography or black hole physics or cosmologies. While a clean definition of path-integral in ordinary non-gravitational QFT can be arranged, such a luxury is not achievable in gravitational systems which suffers from perturbative non-renormalizability, suggesting an effective treatment of the problem tHooft:1974toh ; Deser:1974nb ; Deser:1974cz ; Goroff:1985sz ; Goroff:1985th ; vandeVen:1991gw . Besides this, the choice of contour enabling flat spacetime QFT path-integral convergent is not available. On the other hand, the conformal mode leading to unboundedness of the Euclidean gravitational path-integral is suggestive of working with the gravitational path-integral directly in the Lorentzian signature which, however, is highly oscillatory Gibbons:1978ac . As a result, the path-integral over the geometries becomes a challenging problem which is sensitive to boundary choices and contours in the space of all possible complex metrics.
@@
In a semi-classical approximation, the gravitational path-integral reduces to a summation over contributions coming from various saddle geometries. Typically for given boundary choices, multiple saddle-geometries (which can be complex) appear in path-integral competing with each other. Dominance of one over the other is dictated by the external inputs, where the saddles can loose their dominance as external parameters are varied. This is analogous to phase transitions occurring in non-gravitational settings. Understanding dominance of saddles is therefore crucial in gaining a deeper insight to various phases of the gravitational systems.
@@
However, when the path-integral over metrics is extended to complex space, subtlety arises. Not all such metrics are allowable, where quantum field theories can be meaningfully and sensibly defined Witten:2021nzp . Kontsevich-Segal-Witten (KSW)-allowability criterion offers a diagnostic tool to identify such complex metrics. This criterion directly arises from the requirements of convergence and well-behavedness of the path-integral, bifurcating the space of complex geometries in to allowable and non-allowable ones. A crucial question then arises regarding the allowability of various complex saddles in the path-integral, asking piercing questions regarding the region where they lie. This however is sensitive to the choice of external parameters for example boundary configurations, in the sense that saddle allowability is affected by the boundary choices.
@@
In the context of quantum cosmology, it is seen that different boundary choices at initial time lead to a very different Universe in the future. By demanding a Universe to be devoid of an initial singularity, a Universe starting from zero size (No-boundary Universe) is proposed. In a simple mini-superspace model it is achievable by either fixing initial size of the Universe (Dirichlet BC) or by fixing the initial momentum (Neumann BC) DiTucci:2019bui ; Narain:2021bff ; Lehners:2021jmv ; DiTucci:2020weq ; Narain:2022msz ; Ailiga:2023wzl ; Ailiga:2024mmt . In the former multiple saddles compete for dominance in various regimes given by final size of the Universe. The complex saddles which eventually becomes relevant in the path-integral computed via Picard-Lefschetz methods turn out to have unstable perturbations Feldbrugge:2017kzv ; DiTucci:2018fdg ; Feldbrugge:2017fcc . In the Neumann case, the complex saddles corresponding to Hartle-Hawking (HH) no-boundary geometries, leads to stable perturbations Lehners:2021jmv ; Ailiga:2024wdx . The saddle corresponding to the tunneling proposal is not only sub-dominant but also has unstable perturbations. Despite these appealing features, the HH-saddle however, gives rise to great puzzles: it predicts an empty Universe with least possible number of inflationary e-folds and predicts an overall spatial curvature of the Universe that is significantly larger than the observed one Maldacena:2024uhs .
@@
Recently, it has been noticed that saddle configurations corresponding to wine-glass geometries (first introduced in Lavrelashvili:1988un ), eases this tension Betzios:2024oli ; Betzios:2024zhf ; Betzios:2026rbv , though such saddles may not enjoy the privilege of being dominant in the path-integral as noticed in Lavrelashvili:2026zsw . These are wormhole saddles where the Euclidean wormhole bridges the two asymptotic regions, and are intrinsically non-perturbative configurations following from equation of motion. They are Euclidean wormhole saddles supported by Axion-field, and in situations where they are dominant saddles over others, can be analytically continued to Lorentzian signature, where they lead to cosmological evolution. A particular feature of these wormhole geometries is that they are asymptotically AdS in the far past. These gravitational saddles are noticed to overcome the various issues affecting the Hartle-Hawking no-boundary saddle i.e. supports long lasting inflationary period. Though it is still not well-established how from a single boundary configuration both wineglass saddles and no-boundary instanton can both arise, where the wineglass saddle emerges as a dominant one.
@@
In this short note we check the KSW allowability of particular type of wineglass geometries that are obtained via analytical continuation of Lorentzian deSitter solution. This is important because if such geometries are dominant for some boundary configurations and are KSW-allowed then they offer a favourable proposal over the Hartle-Hawking no-boundary instanton as it overcomes various issues by giving preference to longer inflationary phase. On the other hand if they are not KSW-allowed, which may very well happen, as has been noticed for cases like bouncing cosmologies or typical wormholes geometries, then they lead to unphysical geometries on which sensible QFTs cannot be defined. It is therefore crucial to address this issue. In this short note we do a semi-rigorous study to analyse the KSW-allowability of the wineglass geometries which are analytically continued from Lorentzian dS. These, however, do not cover wineglasses which are obtained via analytic continuation of Euclidean AdS Betzios:2024oli ; Betzios:2024zhf ; Betzios:2026rbv ; Lavrelashvili:2026zsw .
The rest of the paper is organized as follows: In Sec. (II), we briefly review the KSW criterion. Sec. (III), details about the techniques used to check the admissibility complex geometry and particularly applied for No-boundary instanton, while Sec. (IV) deals with the admissibility of wine-glass geometry. Finally, Sec. (V) presents summary of our findings and conclusions of the paper.
@@ IV Allowability of Wine-Glass geometry
After having studied the admissibility of the NB instanton, we now proceed to study another interesting complex geometry - Wineglass wormholes with asymptotically Euclidean AdS in the far past, see figure (1(b)). Although wineglass geometries can be obtained in various ways, however the one we focus on in this paper is obatined via analytic continuation of Lorentzian deSitter. These are obtained in pure gravity. It is crucial to point out that these wineglass-geometries doesn’t arise directly as saddle solutions to the EH action, but rather are artificially constructed via analytic continuation of Lorentzian dS which is a saddle of Einstein-Hilbert gravity.
Recent studies involving wineglass-geometries which are supported by matter (magnetic or axionic fields) have come across saddles which are dominant compared to others. Geometrically these saddles are Euclidean AdS which are analytically continued to lead to Lorentzian dS Jonas:2023ipa ; Betzios:2024oli ; Betzios:2024zhf ; Betzios:2026rbv ; Lavrelashvili:2026zsw . Phenomenologically they lead to a prolonged period of inflation, potentially resolving some of the shortcomings of the no-boundary proposal. While such wineglass saddles are quite interesting, they are technically a bit involved to analyse from the perspective of allowability.
In the present work therefore as a first attempt to analyse the KSW-allowability of wineglass geometries, we limit our attention to pure-gravity and follow the time contour approach presented in Hertog:2011ky . In Hertog:2011ky , utilizing the appropriate time contour, the authors showed that in the context of AdS/CFT, the no-boundary saddle admits a representation of regular Euclidean AdS domain wall that makes a smooth transition to a asymptotically deSitter. The representation discussed demonstrates the realization of an AdS-to-dS transition through a complex geometry. It is emphasized that contributions from the transiting complex geometry regulate the volume divergences in the AdS phase and account for the emergence of late-time classical behavior in the de Sitter phase. Furthermore, the regularized Euclidean AdS action of the contour gives the probability amplitude of Lorentzian deSitter history.
@@
Figure 3: A visualization of wineglass geometry and the corresponding time contour. In fig (a): the light-orange correspond to Euclidean AdS, the gray-region corresponds to complex geometry, is the wormhole joining EAdS to EdS, the blue-region is the section of 44-sphere which is the Euclidean deSitter, the white-region is the Lorentzian deSitter. Fig (b) correspond to the time-contour: the vertical red line correspond to EAdS, the circular region depicts the wormhole with throat, the horizontal black line correspond to section of spherical geometry, while the vertical blue line correspond to Lorenztian dS.
@@
This is a negative signature representation of the geometry of Euclidean anti-deSitter space, with a cosmological constant $\bar{\Lambda}=-\Lambda$. Recent studies Betzios:2024oli ; Betzios:2024zhf ; Betzios:2026rbv have shown that contributions from this contour will have higher weighting in the path-integral leading to long lasting inflationary period. Now applying KSW criterion to the metric along this contour, we find
@@
which clearly shows the violation of KSW condition. This hints atleast in case of pure gravity the time contour realization of the wine-glass geometry is KSW disallowed, while NB saddle is KSW allowed. A better understanding can be gained by analyzing the $q=0$ admissibility condition i.e $\mathcal{R}e(\sqrt{g})>0$, along a complex circular contour (cyan contour in fig. (3)) that glues the Euclidean AdS geometry with the Euclidean sphere. Along the circular contour, $\tau=Re^{i\theta}$, with $\theta\in[-\pi/2,0]$ the geometry given in eq. (7) reads as:
@@
From the plot (3), it is clearly evident to see the violation of $q=0$ admissibility condition for some $\theta$ interval. This supports the argument that wine-glass geometry achieved via analytic continuation in pure gravity with cosmological constant is KSW disallowed. Interesting, in the vanishing throat limit i.e. $R\to 0$ one achieves no-boundary instanton, which is KSW - allowed. Furthermore, for better and conclusive insight, we perform the Ridge criterion analysis. The Ridge criterion for wine-glass is illustrated in the figure (3). The analysis reveals that the wine-glass geometry — moreover, any asymptotically EAdS configuration — violates the KSW criterion. This is because, as illustrated in Fig.(3), the asymptotically EAdS region is surrounded by the disallowed region, implying that there exists no allowable contour connecting an asymptotically EAdS geometry to an asymptotically de Sitter configuration. This hints atleast from the time contour representation point of view, that is the wine-glass geometry obtained via analytic continuation from Lorentzian dS in pure-gravity are KSW disallowed, while NB saddle is KSW allowed.
@@
In this short paper we consider class of geometries which resemble like wine-glass and have recently gained attention in the area of quantum cosmology due to the ability of some of them to have a longer phase of inflation thereby overcoming the issues affecting the Hartle-Hawking no-boundary instanton. These wine-glasses are wormholes which typically connect the Euclidean AdS in the far past to a section of Euclidean $S^{4}$ which glues to Lorentzian deSitter at the equator (see figure 1(b)). While they come under same family as the No-boundary instanton which emerges from wine-glass geometries in the limit of vanishing throat size (pinching scenario), their dominance over no-boundary instanton or vice-versa is yet to be settled Lavrelashvili:2026zsw .
@@
We then apply these tools to the wineglass-geometries. The wine-glass geometries discussed in this paper are achieved via analytic continuation of Lorentzian deSitter solution in pure gravity. These are different from the wineglass-geometries studied in Betzios:2024oli ; Betzios:2024zhf ; Betzios:2026rbv ; Lavrelashvili:2026zsw , where the wormhole is supported by axionic (or magnetic) charges, and are analytic continuation of EAdS to future with Lorentzian dS. The wineglass-geomtries studied in this paper are different from that and have time contour shown in figure (3). In our case geometries with negligible throat size correspond to contours involving sharp turns at the origin of the complex $\tau$-plane, which is singular ($R\to 0$). Non-zero throat size involve contours starting at $\tau_{y}=$-$\infty$, where the geometry is EAdS. It stops at $\tau_{y}=-\tau^{(0)}_{y}<0$, leading to a non-zero throat size. The contour then ventures in complex plane to connect to No-boundary instanton contour starting at $\tau_{x}=-\tau^{(0)}_{x}>0$. The transition from EAdS to EdS happens entirely via complex geometries.
@@
Points lying in the far past correspond to EAdS. This mentioned in eq. (17) shows that all entries of the metric are negative, immediately indicating its KSW disallowability. While this is not conclusive enough, we make use of KSW allowability criterion for $q=0$-form fields. This corresponds to checking ${\rm Re}(\sqrt{g})>0$ for various points on the complex plane, thereby dividing the whole complex-$\tau$ plane into two parts. This is shown in figure (a) of 4 (green region consist of allowable geometries according to requirement ${\rm Re}(\sqrt{g})>0$, while light-orange region correspond to disallowable geometries). This is a bit crude test but gives sufficient hints. It can be clearly seen from figure (a) of 4 that the wormhole (cyan-line) lies partially in allowed and partially disallowed region. This indicates that KSW-allowability following from 0-form field is compromised for this type of wine-glasses.
@@
To strengthen our analysis, we check the KSW-allowability by using ridge criterion. As this test is comparatively more definitive, it offers further assurance of KSW-disallowability of class of wineglass geometries investigated in this paper. This test shows (see figure (b) in 4) that while contours corresponding to Euclidean (black-line) and Lorentzian (blue-line) deSitter are lying in KSW allowed region, the contour (red-line) corresponding to EAdS lies in KSW-disallowed region. The contour corresponding to the wormhole (cyan-line) lies partially in allowed and partially in disallowed region. This indicates that contour corresponding to wineglass focued on in this paper are KSW-disallowed.
@@
One must however understand that KSW-allowability of geometries depend on quite a few things and can get affected depending on boundary choices, presence of additional fields and their interactions. The present study doesn’t take those into consideration. The conclusions obtained in this paper cannot be naively generalized to all possible kind of wineglass-geometries, while it may however does provide some hint of their allowability fate which needs to be further confirmed by a proper investigation. The allowability test conducted here are for wineglass geometries in pure-gravity, which are obtained via analytic continuation of deSitter. These are seen to KSW-disallowed. A more rigorous study involving extremal curves along with presence of additional interacting fields leading to more generic types of wineglass-geometries Betzios:2024oli ; Betzios:2024zhf ; Betzios:2026rbv ; Lavrelashvili:2026zsw is still missing in the current work and we hope to address those in the future investigations.